The charge $C, for hiring a hall for an event is C = 150 + 2N, where N stands for the number of people at the event. Find:
a) the charge when 225 people are at the event
b) the number of people at the event when the charge is $394

Answer by jim_thompson5910(35256)

Part (a)

Plug in N = 225 to find C

C = 150 + 2N
C = 150 + 2*225
C = 150 + 450
C = 600

It will cost $600
======================================================================
Part (b)

Plug in C = 394 and solve for (or isolate) N. The idea is to undo PEMDAS.

C = 150 + 2N
394 = 150 + 2N
150 + 2N = 394
2N + 150 = 394
2N + 150 - 150 = 394 - 150 subtract 150 from both sides
2N = 244
2N/2 = 244/2 divide both sides by 2
N = 122

There are 122 people at the event.